# Spatially-Grounded Document Retrieval via Patch-to-Region Relevance Propagation

### v1

@@ 5 Implementation: Snappy System
We implement our approach in Snappy, an open-source document retrieval system available at https://github.com/athrael-soju/Snappy. The system demonstrates the practical applicability of spatially-grounded hybrid retrieval.
@@ 5.1 Architecture Overview
Figure 1 illustrates the Snappy system architecture, showing the parallel document indexing pipeline and query processing pipeline that converge at region-level filtering.
Figure 1: Snappy system architecture. The document indexing pipeline processes uploads through parallel OCR and ColPali embedding branches, storing regions in DuckDB and patch vectors in Qdrant. The query processing pipeline retrieves top-K candidates, generates interpretability maps from patch-level attention, and filters OCR regions by relevance to produce answers with spatial context.
Snappy pairs a FastAPI backend with a ColPali embedding service, DeepSeek OCR for visual grounding, DuckDB analytics for metadata storage, and a Next.js frontend. The system delivers hybrid vision+text retrieval over PDFs, where each page is rasterized, embedded as multivectors, and stored alongside images and OCR-extracted text in Qdrant.
The indexing pipeline processes documents as follows: (1) render each page as an image using Poppler, (2) extract patch embeddings via the ColPali service, (3) optionally extract text regions with bounding boxes via DeepSeek OCR with visual grounding enabled, (4) store patch embeddings as multivectors and OCR metadata in Qdrant with page-level pooled embeddings for efficient retrieval, and (5) persist images to MinIO object storage for visualization.
@@
At query time: (1) encode the query via ColPali’s text encoder, (2) retrieve top-$K$ candidate pages via ANN search on pooled embeddings in Qdrant, (3) compute full patch-level similarity for candidates using the stored multivectors, (4) propagate scores to OCR regions via IoU-weighted aggregation, and (5) return ranked regions with text content, bounding boxes, and optional visual highlighting.
5.2 Component Integration
@@
The ColPali embedding service runs as a dedicated container exposing patch-level embeddings over HTTP, supporting both GPU and CPU deployment. Unlike standard ColPali deployments that return only page-level scores, Snappy’s integration extracts the full $(n\times G^{2})$ similarity matrix stored as multivectors in Qdrant for downstream region scoring.
DeepSeek OCR provides text extraction with visual grounding capabilities. When enabled, the OCR service extracts markdown-formatted text alongside bounding box coordinates for each text region. OCR results are stored in DuckDB for SQL-based querying.
Qdrant stores both page-level pooled embeddings (for Stage 1 candidate retrieval) and full patch embeddings as multivectors (for Stage 2 region scoring).
6 Planned Evaluation
Empirical evaluation is ongoing using BBox-DocVQA (Yu et al., 2025) (3.6K documents with 32K question-answer pairs annotated with evidence bounding boxes) and ViDoRe for page-level baselines. We will measure retrieval quality (NDCG@k, Precision@k), spatial grounding accuracy (IoU@0.5, IoU@0.7), and downstream RAG answer quality (ANLS). Baselines include ColPali page-level retrieval, BM25 on OCR text, and dense retrieval on OCR chunks. Ablations will compare aggregation strategies and threshold sensitivity. Results will be reported in a subsequent revision.
@@ 7.1 Limitations
Patch Resolution Bound. As analyzed in Section 4, small regions suffer from limited localization precision due to fixed patch granularity. For ColPali’s 14-pixel patches, regions smaller than approximately $35\times 35$ pixels achieve less than 50% precision. Future work could explore multi-scale patch embeddings or dynamic resolution based on document density.
OCR Dependency. Region quality depends on OCR accuracy and segmentation. Poorly segmented regions (merged paragraphs, split tables) degrade retrieval quality regardless of the VLM’s spatial attention accuracy. Layout-aware OCR systems like DeepSeek OCR with visual grounding mitigate this, but OCR errors propagate to retrieval.
@@
Two-Stage Information Loss. The mean-pooling in Stage 1 discards spatial information, potentially missing pages where relevant content is concentrated in small regions. Relevant pages could be filtered out before Stage 2 region scoring. Alternative pooling strategies (max pooling, attention-weighted pooling) may mitigate this.
Theoretical Assumptions. The performance bounds in Section 4.3 assume that ColPali’s patch scores correlate positively with region relevance. While this is plausible given ColPali’s strong page-level retrieval performance, the degree of correlation at the region level remains to be empirically validated.

### v2

@@ 5 Implementation: Snappy System
We implement our approach in Snappy, an open-source document retrieval system available at https://github.com/athrael-soju/Snappy.
@@ 5.1 Architecture Overview
Snappy implements a two-stage retrieval pipeline over PDFs. Each page is rendered as an image, embedded via ColPali to produce patch-level multivectors, and optionally processed by layout-aware OCR to extract text regions with bounding boxes. Both patch embeddings and OCR metadata are stored in a vector database for retrieval.
@@
At query time, the system: (1) encodes the query via ColPali’s text encoder, (2) retrieves top-$K$ candidate pages via ANN search on pooled embeddings, (3) computes full patch-level similarity for candidates, (4) propagates scores to OCR regions via IoU-weighted aggregation, and (5) returns ranked regions with text content and bounding boxes.
@@
Unlike standard ColPali deployments that return only page-level scores, Snappy extracts the full $(n\times G^{2})$ similarity matrix for downstream region scoring. The vector database stores both page-level pooled embeddings (Stage 1) and full patch multivectors (Stage 2). The overall architecture is illustrated in Figure˜1.
Figure 1: Snappy system architecture. The indexing pipeline processes documents through parallel OCR and embedding branches. The query pipeline retrieves candidates via ANN search, computes patch-level similarities, and filters OCR regions by relevance.
6 Empirical Evaluation
We evaluate our approach on BBox-DocVQA, measuring spatial grounding accuracy and token efficiency. Our experiments address three questions: (1) How accurately does patch-to-region relevance propagation localize relevant content? (2) How does performance vary across document categories? (3) What token savings does the approach achieve?
Evaluation Scope. Our evaluation targets the core contribution: region-level localization within a page. Given a page containing the answer, can our approach identify the specific region? This complements page-level retrieval benchmarks like ViDoRe (Faysse et al., 2025), which measure whether the correct page is retrieved. We use IoU-based spatial grounding metrics to directly measure localization accuracy against ground-truth evidence bounding boxes. We evaluate within-page localization (Stage 2); page-level retrieval accuracy (Stage 1) is measured by existing ViDoRe benchmarks and is outside our scope.
6.1 Experimental Setup
Dataset. BBox-DocVQA (Yu et al., 2025) provides question-answer pairs across documents from arXiv categories including computer science (cs), economics (econ), electrical engineering (eess), mathematics (math), physics, quantitative biology (q-bio), quantitative finance (q-fin), and statistics (stat). Each QA pair includes ground-truth bounding boxes marking the evidence region containing the answer. We evaluate on 1,619 samples across eight categories (1,623 total samples minus 4 that consistently failed during OCR processing and are excluded from all reported metrics).
Models. We evaluate three ColPali-family models: ColQwen3-8B (8B parameters) and ColQwen3-4B (4B parameters) representing state-of-the-art accuracy at different scales, and ColModernVBERT (250M parameters) representing the efficiency frontier.
Configuration. We use DeepSeek-OCR with visual grounding in markdown mode, selected for its layout-aware region segmentation that groups semantically coherent blocks (paragraphs, table cells, captions) rather than individual words or lines. For region scoring, we apply 50th-percentile thresholding (denoted P50) with max token aggregation and max region scoring. The P50 threshold represents a principled midpoint validated by our ablation study (Section˜6.8): 25th-percentile thresholds (P25) are unnecessarily inclusive (all regions selected), while 75th-percentile thresholds (P75) require careful tuning of minimum patch overlap to avoid accuracy degradation.
Metrics. For predicted region $B_{p}$ and ground-truth bounding box $B_{g}$, we compute:
$\text{IoU}(B_{p},B_{g})=\frac{|B_{p}\cap B_{g}|}{|B_{p}\cup B_{g}|}$ (13)
We report:
  • Mean IoU: Average IoU between selected regions and ground-truth bounding boxes
[…]
  • Token Savings: Percentage reduction in context tokens, computed as $(T_{\text{baseline}}-T_{\text{method}})/T_{\text{baseline}}$, compared to (a) all OCR regions and (b) full image tokens
6.2 Main Results
Table 3: Spatial grounding accuracy on BBox-DocVQA. Hit rates indicate the fraction of samples achieving IoU at or above the specified threshold.
Model N Mean IoU IoU@0.25 IoU@0.5 IoU@0.7
ColModernVBERT (P50) 1619 0.480 72.2% 45.5% 27.9%
ColQwen3-4B (P50) 1619 0.569 84.4% 59.7% 35.8%
ColQwen3-8B (P50) 1619 0.569 83.9% 59.8% 36.1%
Figure 2: Hit rate comparison across IoU thresholds. ColQwen3-8B and ColQwen3-4B achieve nearly identical performance, while both consistently outperform ColModernVBERT, with the gap widening at stricter thresholds.
ColPali’s patch attention localizes relevant regions without additional training (Table˜3). ColModernVBERT (250M parameters) achieves 45.5% hit rate at IoU@0.5 with the highest token efficiency (38.0% savings vs OCR, 58.5% vs full image). Scaling to ColQwen3-4B improves localization substantially (59.7% at IoU@0.5), but further scaling to ColQwen3-8B yields negligible gains (59.8%), demonstrating diminishing returns from parameter scaling within the same architecture (Figure˜2).
6.3 Category Analysis
Table 4: Performance by document category across all three models. ColQwen3-8B and ColQwen3-4B achieve nearly identical accuracy, while ColModernVBERT shows consistent degradation. Mathematics and economics documents show notably lower accuracy across all models, likely due to denser tabular content and smaller region sizes where patch quantization effects are more pronounced.
ColModernVBERT ColQwen3-4B ColQwen3-8B
Category N Mean IoU IoU@0.5 Mean IoU IoU@0.5 Mean IoU IoU@0.5
cs 216 0.527 49.5% 0.697 75.5% 0.689 74.5%
eess 196 0.451 42.9% 0.656 78.1% 0.660 79.1%
q-bio 176 0.575 59.1% 0.616 66.5% 0.616 66.5%
physics 213 0.521 53.5% 0.579 63.8% 0.590 65.7%
stat 200 0.510 50.0% 0.559 57.0% 0.559 57.0%
q-fin 216 0.459 45.8% 0.543 59.7% 0.545 60.2%
econ 214 0.427 36.0% 0.513 46.7% 0.509 46.3%
math 188 0.370 27.1% 0.382 28.7% 0.378 27.7%
Overall 1619 0.480 45.5% 0.569 59.7% 0.569 59.8%
Performance varies substantially across document categories (Table˜4). Computer science and electrical engineering documents achieve substantially higher localization accuracy with ColQwen3-4B (Mean IoU $\geq$ 0.65) than economics (0.513) and mathematics (0.382) documents, with physics (0.579) and quantitative biology (0.616) falling in between. ColModernVBERT shows a similar pattern but with uniformly lower scores; notably, quantitative biology achieves the highest accuracy (0.575) for the smaller model, while computer science drops from 0.697 (ColQwen3-4B) to 0.527 (ColModernVBERT). This performance gap likely reflects differences in layout complexity: economics papers frequently contain dense tables with many small cells, while mathematics papers feature equations and formulas that may span multiple regions. These smaller regions suffer disproportionately from patch quantization effects, consistent with the theoretical area efficiency bounds in Section˜4.
6.4 Token Efficiency
Table 5: Token efficiency comparison. Our hybrid approach with P50 filtering achieves substantial savings over both returning all OCR regions and using full image tokens.
Method Total Tokens vs All OCR vs Full Image
Full Image (baseline) 4,003,039
All OCR Regions 2,678,723 33.1%
ColModernVBERT (P50) 1,661,684 38.0% 58.5%
ColQwen3-4B (P50) 1,908,329 28.8% 52.3%
ColQwen3-8B (P50) 1,974,263 26.3% 50.7%
Figure 3: Token usage comparison across methods. All hybrid approaches achieve substantial reductions compared to full image (4.00M) and all OCR (2.68M) baselines, with ColModernVBERT offering the greatest efficiency at 1.66M tokens. ColQwen3-8B (1.97M) and ColQwen3-4B (1.91M) achieve similar token savings.
Figure 4: Multi-dimensional comparison of ColQwen3-8B, ColQwen3-4B, and ColModernVBERT. ColQwen3-8B and ColQwen3-4B achieve nearly identical accuracy metrics, while ColModernVBERT offers the best token savings. Mean IoU is scaled to percentage for visual consistency.
We measure context tokens that would be passed to a downstream LLM. For full image retrieval, we estimate tokens using Claude’s approximation: images are resized to fit within $1568\times 1568$ pixels while maintaining aspect ratio, then token count is computed as $\lfloor(w\times h)/750\rfloor$. For OCR-based methods, we tokenize extracted text using tiktoken’s cl100k_base encoding (GPT-4 tokenizer, used as approximation). Token counts are summed across all 1,619 evaluation samples. The hybrid approach with P50 filtering reduces context tokens by 28.8% compared to returning all OCR regions and by 52.3% compared to full-page image tokens (Tables˜5 and 3). These savings directly reduce LLM inference costs and context window usage, consistent with the theoretical bounds in Corollary˜2. The accuracy-efficiency tradeoff in Figure˜4 confirms that ColQwen3-8B and ColQwen3-4B achieve nearly identical accuracy, while ColModernVBERT offers the best token savings.
6.5 Model Architecture and Training Data Effects
The substantial performance gap between ColQwen3-4B and ColModernVBERT (14.2 percentage points at IoU@0.5) invites analysis of how model architecture and training data affect patch-to-region relevance propagation. Notably, ColQwen3-8B achieves nearly identical performance to ColQwen3-4B (59.8% vs 59.7% at IoU@0.5), suggesting diminishing returns from parameter scaling within the same architecture family.
[…]
Attention Sharpness and Localization. For region-level localization, however, attention sharpness matters more than global context. Localization accuracy depends on how tightly the model’s attention concentrates on query-relevant patches. ColQwen3-4B’s larger capacity produces sharper attention distributions, while its multi-scale features and explicit 2D spatial encoding help attention patterns respect spatial boundaries. These factors explain the 14.2 percentage point IoU@0.5 gap.
Training Data Composition. The models differ substantially in both pre-training scale and fine-tuning data:
  • ColQwen3-4B: Built on Qwen3-VL pre-trained on $\sim$1.4 trillion multimodal tokens (Qwen Team, 2025; Wang et al., 2024) including extensive OCR data, document formats, and VQA datasets. Fine-tuned on VDR (Visual Document Retrieval), ViDoRe-ColPali-Training, and VisRAG-Ret-Train (Shi et al., 2024), with 63% academic datasets (DocVQA, InfoVQA, TAT-DQA) and 37% synthetic query-document pairs.
  • ColModernVBERT: Pre-trained on $\sim$2 billion text tokens with modality alignment on The Cauldron 2 and Docmatix (21% OCR/document content), then contrastively trained on 118k document-query pairs with 300k text-only pairs using a 2:1 text-to-image ratio.
[…]
Scaling Within Architecture Family. The near-identical performance of ColQwen3-8B and ColQwen3-4B (Mean IoU 0.569 for both; 59.8% vs 59.7% at IoU@0.5) reveals that doubling parameters within the same architecture provides negligible localization gains. This suggests that ColQwen3’s localization accuracy is bounded by factors other than model capacity, likely patch resolution and attention pattern characteristics rather than representational capacity. The practical implication is clear: ColQwen3-4B offers the same localization quality at lower computational cost.
Category-Specific Performance Gaps. Table˜6 reveals that the performance gap between ColQwen3-4B and ColModernVBERT varies systematically across document categories.
Table 6: Performance gap analysis between ColQwen3-4B and ColModernVBERT by category. ColQwen3-8B results (not shown) are within 0.01 Mean IoU of ColQwen3-4B across all categories. The gap narrows substantially for mathematics documents, suggesting both model families approach fundamental precision limits for small-region localization.
Category ColQwen3-4B ColModernVBERT $\Delta$ Mean IoU $\Delta$ IoU@0.5
cs 0.697 0.527 -0.170 -26.0 $p$
eess 0.656 0.451 -0.205 -35.2 $p$
q-bio 0.616 0.575 -0.041 -7.4 $p$
physics 0.579 0.521 -0.058 -10.3 $p$
stat 0.559 0.510 -0.049 -7.0 $p$
q-fin 0.543 0.459 -0.084 -13.9 $p$
econ 0.513 0.427 -0.086 -10.7 $p$
math 0.382 0.370 -0.012 -1.6 $p$
Overall 0.569 0.480 -0.089 -14.2 $p$
Figure 5: IoU distribution by document category for ColQwen3-8B, ColQwen3-4B, and ColModernVBERT. Box plots show median, quartiles, and outliers. Categories ordered by mean IoU (highest to lowest). Mathematics documents show consistently low performance across all models, while computer science achieves both higher mean and greater spread.
The IoU distribution across categories (Figure˜5) reveals that the smallest performance gap occurs in mathematics documents ($\Delta=0.012$ Mean IoU), where both models achieve their lowest absolute performance. This convergence suggests that mathematics documents present challenges that affect both models similarly, whether due to training data composition, difficulty parsing equations and dense notation, or other factors. The precise cause remains an open question.
Conversely, the largest gaps appear in computer science ($\Delta=0.170$) and electrical engineering ($\Delta=0.205$) documents, where ColQwen3-4B’s superior attention discrimination yields substantial improvements. These categories feature larger, well-separated regions (figures, code blocks, prose paragraphs) where model capacity translates directly to localization accuracy.
Implications for Model Selection. This analysis suggests three empirically distinct regimes:
  1. 1.
    Architecture-sensitive documents: For documents with large, well-separated regions (CS, EESS, Q-Bio), ColQwen3 models substantially outperform ColModernVBERT.
  2. 2.
    Scale-insensitive within architecture: ColQwen3-8B and ColQwen3-4B achieve nearly identical performance, indicating that doubling parameters within the same architecture yields negligible localization gains.
  3. 3.
    Uniformly challenging documents: For mathematics and dense tabular content, all models converge to similar (lower) performance, suggesting fundamental precision limits.
[…]
6.6 Error Analysis
To understand localization failures, we analyze the 40.3% of samples (652/1,619) where the top-ranked region fails to achieve IoU $\geq$ 0.5 with the ground-truth bounding box (the complement of the 59.7% hit rate in Table˜3). Among these failures, we identify 487 samples where no selected region achieves adequate overlap, which we analyze by root cause. The error decomposition in Table˜7 reveals two failure modes: (1) OCR ceiling failures, where OCR segmentation did not produce any region with IoU $\geq$ 0.5 to the ground truth, and (2) selection failures, where OCR successfully segmented the target region but our relevance propagation selected a different region.
Table 7: Error analysis for ColQwen3-4B failures at IoU@0.5. The dominant failure mode (85%) is OCR segmentation ceiling: our approach cannot select regions that do not exist. Only 15% of failures are attributable to the relevance propagation mechanism.
Failure Mode Count % of Failures Description
OCR ceiling 414 85.0% No OCR region overlaps GT sufficiently
Selection error 73 15.0% OCR found target; wrong region selected
Total failures 487 100%
OCR Ceiling Dominates. The majority of failures (85%) occur because the OCR system did not segment a region that adequately overlaps the ground-truth bounding box. This represents a fundamental ceiling: our relevance propagation cannot select regions that do not exist. This finding validates our architectural choice: the patch-to-region propagation mechanism works well when given appropriate candidates.
Selection Failures by Category. Among the 73 selection failures (where OCR found the target but we selected incorrectly), the distribution varies by category: q-bio (9.7% selection error rate), physics and eess (6.1%), cs (2.3%), and econ (0.9%). Categories with higher selection error rates may feature more semantically ambiguous layouts where multiple regions could plausibly contain the answer.
Partial Overlap in Failures. Even among samples that fail at IoU@0.5, 48% achieve IoU $\geq$ 0.25, indicating substantial partial overlap. The mean IoU of failed samples is 0.250, suggesting that failures often represent near-misses rather than complete mislocalization. This has practical implications: for applications tolerant of approximate localization, the effective accuracy is higher than the IoU@0.5 metric suggests.
Implications. This error analysis suggests two directions for improvement: (1) better OCR segmentation to raise the ceiling, potentially through region merging or multi-scale segmentation, and (2) improved attention sharpness for the 15% of cases where selection fails despite adequate OCR coverage. We discuss these limitations and future directions further in Section˜7.
6.7 Qualitative Examples
(a) Query asks which letter receives two arrows in a figure. Patch attention concentrates on the diagram region; predicted box (magenta) substantially overlaps ground truth (green).
(b) Query asks about table cells with Roman numerals. Attention focuses on the table structure; multiple OCR regions (blue) are scored, with the table region selected.
Figure 6: Patch attention heatmaps with region overlays. Red intensity indicates patch-level similarity scores from ColQwen3-4B. Green: ground-truth bounding box. Blue: OCR-detected regions. Magenta: predicted (selected) regions. The heatmaps demonstrate that patch attention spatially concentrates on query-relevant content, enabling accurate region selection when OCR segmentation aligns with semantic boundaries.
The patch attention heatmaps (Figure˜6) illustrate our relevance propagation mechanism on two representative samples. The heatmap overlay (red intensity) shows ColQwen3-4B’s per-patch similarity scores, revealing where the model attends when processing the query. Ground-truth bounding boxes (green), OCR-detected regions (blue), and our predicted regions (magenta) demonstrate the spatial alignment between patch attention and region selection. The left example (Figure˜6(a)) shows a successful localization where the query references a specific figure. Patch attention strongly activates over the diagram containing the answer, and our IoU-weighted aggregation correctly selects the corresponding OCR region. The right example (Figure˜6(b)) demonstrates table localization, where attention distributes across the table structure and our approach selects the relevant table region despite multiple OCR candidates on the page.
These visualizations illustrate the core mechanism: ColPali’s emergent patch attention, trained only for page-level retrieval, provides sufficient spatial signal to discriminate between OCR regions when propagated through our IoU-weighted scoring.
6.8 Threshold Sensitivity Analysis
We analyze sensitivity to three hyperparameters: percentile threshold (P25, P75), region scoring method (max, weighted average), and minimum patch overlap (0.1, 0.25, 0.5). We evaluate on an 80-sample stratified subset (10 per category) across 12 configurations to enable systematic comparison.
Parameter-Insensitive Regime at P25. All 6 configurations using P25 thresholding produce identical results (Mean IoU 0.605, IoU@0.5 68.8%), regardless of region scoring or minimum overlap. This demonstrates that at sufficiently inclusive thresholds, the filtering mechanism admits nearly all regions, rendering other parameters irrelevant.
Parameter Sensitivity at P75. Higher percentile thresholds introduce meaningful parameter sensitivity, as shown in Table˜8.
Table 8: Ablation study results. All P25 configurations produce identical results regardless of region scoring or minimum overlap. At P75, minimum overlap of 0.1 causes accuracy degradation, with weighted average scoring showing greater robustness than max scoring.
Region Scoring Min Overlap Mean IoU IoU@0.25 IoU@0.5 $\Delta$ IoU@0.5
P25 (all 6 configurations identical)
any any 0.605 91.2% 68.8%
P75 configurations
max 0.1 0.566 86.2% 63.7% -5.1 $p$
max 0.25 0.605 91.2% 68.8% 0.0 pp
max 0.5 0.605 91.2% 68.8% 0.0 pp
weighted_avg 0.1 0.587 88.8% 66.2% -2.6 $p$
weighted_avg 0.25 0.605 91.2% 68.8% 0.0 pp
weighted_avg 0.5 0.605 91.2% 68.8% 0.0 pp
Minimum Overlap Dominates at P75. The only configurations showing accuracy degradation are those with minimum patch overlap of 0.1. At this threshold, regions are scored based on patches with minimal spatial intersection, introducing noise. Increasing minimum overlap to 0.25 or 0.5 fully recovers baseline accuracy, suggesting that requiring moderate spatial correspondence between patches and regions is necessary for reliable scoring.
Region Scoring Robustness. At low minimum overlap (0.1), weighted average scoring (Mean IoU 0.587) outperforms max scoring (0.566), demonstrating greater robustness to noisy patch contributions. At higher overlap thresholds, both methods produce identical results, as patch-region alignment is sufficient for either aggregation strategy.
Token Aggregation Invariance. We additionally validated that mean and sum token aggregation produce identical per-sample IoU scores on a subset of configurations, confirming that the relative magnitude of token contributions does not affect region ranking. This suggests that patch-level attention patterns, rather than token-specific weighting, drive localization accuracy.
Practical Implications. These findings validate our default configuration (P50, IoU-weighted aggregation) as a robust middle ground. Lower thresholds (P25) are unnecessarily inclusive, while higher thresholds (P75) require careful tuning of minimum overlap. For practitioners, the key insight is that minimum patch overlap should be set to at least 0.25 when using stricter percentile thresholds.
@@ 7.1 Limitations
Patch Resolution Bound. As analyzed in Section˜4, small regions suffer from limited area efficiency due to fixed patch granularity. For ColPali’s 14-pixel patches, regions smaller than approximately $35\times 35$ pixels achieve less than 50% area efficiency. Future work could explore multi-scale patch embeddings or dynamic resolution based on document density.
OCR Dependency. Region quality depends on OCR accuracy and segmentation. Poorly segmented regions (merged paragraphs, split tables) degrade retrieval quality regardless of the retrieval model’s spatial attention accuracy. Layout-aware OCR systems with visual grounding mitigate this, but OCR errors propagate to retrieval. Note that OCR segmentation quality affects region boundaries but is orthogonal to our relevance scoring mechanism; our contribution is patch-to-region propagation, not OCR quality.
@@
Category-Specific Performance. Empirical evaluation reveals substantial performance variation across document types. Mathematics documents achieve only 28.7% hit rate at IoU@0.5 compared to 78.1% for electrical engineering documents. This gap likely reflects multiple factors: training data composition favoring certain document types, difficulty parsing equations and dense mathematical notation, and potential patch quantization effects for small regions. Applications targeting specific document types should consider this variation when setting performance expectations.
Within-Document Variance. Variance decomposition reveals that 72% of IoU variance is within-document (same document, different questions) versus 28% between documents. This suggests that question-specific semantic factors, rather than inherent document difficulty, dominate localization performance. A document may yield high IoU for one question and low IoU for another, depending on how well the query aligns with the model’s learned attention patterns. This finding implies that per-category performance statistics represent averages over highly variable per-question outcomes.